OpenStudy (bloomlocke367):

Tom was asked how old he was. He replied "In 2 years I will be twice as old as I was five years ago.' How old is he ?

OpenStudy (anonymous):

You can't figure it out without how old he was 5 years ago...

OpenStudy (bloomlocke367):

sure you can.

OpenStudy (anonymous):

>:C

OpenStudy (bloomlocke367):

would you like an explanation @AnswerMyQuestions

OpenStudy (bloomlocke367):

@jigglypuff314 @iGreen

OpenStudy (anonymous):

Yes. :C

OpenStudy (bloomlocke367):

okay, i'm gonna message you.. I don't wanna spoil it for someone else..

OpenStudy (anonymous):

xd

OpenStudy (anonymous):

And I'll post the whole thing you message me onto here! Muahahahahaha! C:< jk

OpenStudy (igreen):

x + 2 = 2x - 5 I think that's correct..hold on.

OpenStudy (bloomlocke367):

close...

OpenStudy (igreen):

No..

OpenStudy (bloomlocke367):

do you not know either?

OpenStudy (igreen):

I'm still thinking :P

OpenStudy (bloomlocke367):

ok. take your time :)

OpenStudy (igreen):

I got 12..

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

That's what she sent me...

OpenStudy (igreen):

12 - 5 = 7 12 + 2 = 14 14 = half of 7

OpenStudy (anonymous):

x+2=2(x-5) x+2=2x-10 2=x-10 x=12

OpenStudy (igreen):

:P

OpenStudy (bloomlocke367):

good job :)